# Some topological genera and Jacobi forms

### v2

@@ Abstract.
We revisit and elucidate the $\widehat{A}$-genus, Hirzebruch’s $L$-genus and Witten’s $W$-genus, cobordism invariants of special classes of manifolds. After slight modification, involving Hecke’s trick, we find that the $\widehat{A}$-genus and $L$-genus arise directly from Jacobi’s theta function. In this way, for every $k\geq 0,$ we obtain exact formulas for the quasimodular expressions of $\widehat{A}_{k}$ and $L_{k}$ as “traces” of partition Eisenstein series
$\widehat{\mathcal{A}}_{k}(\tau)=\operatorname{Tr}_{k}(\phi_{\widehat{A}};\tau)\ \ \ \ \ \ {\text{\rm and}}\ \ \ \ \ \ \mathcal{L}_{k}(\tau)=\operatorname{Tr}_{k}(\phi_{L};\tau).$
Surprisingly, Ramanujan defined twists of the $\widehat{\mathcal{A}}_{k}(\tau)$ in his “lost notebook” in his study of derivatives of theta functions, decades before Borel and Hirzebruch rediscovered them in the context of spin manifolds. In addition, we show that the nonholomorphic $G_{2}^{\star}$-completion of the characteristic series of the Witten genus is the Jacobi theta function avatar of the $\widehat{A}$-genus.
@@ 1. Introduction and Statement of Results
This example is historically significant because of its role in the discovery of the Atiyah-Singer index theorem (for example, see Hitchin’s expository article [4]). Atiyah and Singer discovered and employed their index theorem to explain the mysterious integrality of the $\widehat{A}$-genera. To compute these values, they implemented (1.1) with
@@
We prove that $\widehat{A}(p_{1},p_{2},\dots;t)$, after minor modification, is essentially a Jacobi form (see Chapter 2 of [5] or [6]) on $\mathbb{C}\times\mathbb{H}.$ To make this precise, we recall the celebrated Jacobi theta function (see [7, 6])
@@
where $u:=e^{2\pi iz}$ and $q:=e^{2\pi i\tau}.$ This function is a Jacobi form for $\mathrm{SL}_{2}(\mathbb{Z})$ of weight $1/2$ and index $1/2$. We work instead with a slightly modified version of this function Readers familiar with [8] should be aware that $\widetilde{\Theta}(z;\tau)$ here is slightly different from the one in that paper. Namely, in terms of Dedekind’s eta-function $\eta(\tau):=q^{\frac{1}{24}}\prod_{n=1}^{\infty}(1-q^{n}),$ it will be convenient for us to employ
@@
Remark.
The infinite product in (1.5) is taken over relatively prime pairs of integers $(m,n)$ instead of $i=1,2,\dots,$ as in (1.1). This modification does not lose any information. In fact, this reformulation will allow us to compute $\widehat{A}$-genera (and also $L$-genera) as quasimodular forms (see Theorem 1.4). Finally, we note that the dependence on $s>0$ in the index set $X_{\tau}(s)$ is required, as we view these series as analytic functions, and the introduction of $s$ guarantees convergence.
@@
As a consequence of both Theorems 1.1 and 1.2, we obtain quasimodular representations of the $\widehat{A}$-genera and $L$-genera. These forms are given as traces of “partition Eisenstein series,” which are studied in [8, 12, 13]. To define them, recall that a partition of a non-negative integer $k$ (see [14] for background on partitions) is any nonincreasing sequence of positive integers
@@
$\widehat{A}_{3}(p_{1},p_{2},\dots)$ $=\frac{1}{967680}(-16p_{3}+44p_{1}p_{2}-31p_{1}^{4}),$
@@
(1.21) $\operatorname{Tr}_{k}^{(E)}(\phi;\tau):=\sum_{\lambda\vdash k}\phi(\lambda)E_{\lambda}(\tau).$
@@ Acknowledgements
The authors thank Jack Morava for helpful conversations and encouragement. They also thank John Shareshian and Richard Stanley for bringing our attention to Hitchin’s article [4], and they thank Wei-Lun Tsai for spotting typographical errors in an earlier version of this paper. The third author thanks the Thomas Jefferson Fund and the NSF (DMS-2002265 and DMS-2055118).
@@ 2.1. Jacobi forms and elliptic functions
Proof of Proposition 2.1.
Claim (1) follows as an easy application of the Jacobi Triple Product formula (see Theorem 2.8 of [14]), which allows us to write
@@
The second claim follows immediately from (1) and (1.4). ∎
@@
Proof.
The $\sigma$-function has a $q$-series expansion (see Theorem I.6.6.4 of [17]) given by
@@
Thus $\sigma$ is also related to the modified theta function $\Theta(z;\tau)$, defined by (2.1), as claimed. ∎
@@
Proof.
The standard application of “Hecke’s trick” (for example, see p. 84 of [5]), to force convergence of the weight $2$ Eisenstein series, gives the formula
@@
The lemma follows since $\zeta(s)$ is continuous at $2$. ∎
@@
In this section we prove Theorems 1.4 and 1.5, which express the quasimodular representations of the $\widehat{A}$-genera and $L$-genera as traces of partition Eisenstein series. To obtain these results, we make use of exponential generating functions that arose in the previous section. These generating functions can be reformulated as traces of partition Eisenstein series using special identities within the framework of Pòlya’s theory of cycle index polynomials.
@@ 3.1. Pòlya’s cycle index polynomials
The proof of claim (2) is a little more involved. Beginning with (1.17) and Theorem 1.4, we apply the correspondence $G_{2j}\longleftrightarrow B_{2j}s_{j}/2j$ to as follows

### v3

@@ Abstract.
We revisit and elucidate the $\widehat{A}$-genus, Hirzebruch’s $L$-genus and Witten’s $W$-genus, cobordism invariants of special classes of manifolds. After slight modification, involving Hecke’s trick, we find that the $\widehat{A}$-genus and $L$-genus arise directly from Jacobi’s theta function. For every $k\geq 0,$ we obtain exact formulas for the quasimodular expressions of $\widehat{A}_{k}$ and $L_{k}$ as “traces” of partition Eisenstein series
$\widehat{\mathcal{A}}_{k}(\tau)=\operatorname{Tr}_{k}(\phi_{\widehat{A}};\tau)\ \ \ \ \ \ {\text{\rm and}}\ \ \ \ \ \ \mathcal{L}_{k}(\tau)=\operatorname{Tr}_{k}(\phi_{L};\tau),$
which are easily converted to the original topological expressions. Surprisingly, Ramanujan defined twists of the $\widehat{\mathcal{A}}_{k}(\tau)$ in his “lost notebook” in his study of derivatives of theta functions, decades before Borel and Hirzebruch rediscovered them in the context of spin manifolds. In addition, we show that the nonholomorphic $G_{2}^{\star}$-completion of the characteristic series of the Witten genus is the Jacobi theta function avatar of the $\widehat{A}$-genus.
@@ 1. Introduction and Statement of Results
This example is historically significant because of its role in the discovery of the Atiyah-Singer index theorem (for example, see Hitchin’s expository article [4]). Atiyah and Singer discovered and employed their index theorem to explain the mysterious integrality of the values of the $\widehat{A}$-genus. To compute these values, they implemented (1.1) with
@@
We prove that $\widehat{A}(p_{1},p_{2},\dots;t)$ (and the other genera in this paper), after minor modification, is essentially a Jacobi form (see Chapter 2 of [5] or [6]) on $\mathbb{C}\times\mathbb{H}.$ This connection opens the door to new avenues of research, where the theory of modular forms can be brought to bear on the number theoretic properties of these topological invariants that arise as Fourier coefficients. Namely, it is natural to expect that this work will lead to new results about the asymptotic and congruence properties of topological genera.
To make this connection precise, we recall the celebrated Jacobi theta function (see [7, 6])
@@
where $u:=e^{2\pi iz}$ and $q:=e^{2\pi i\tau}.$ This function is a Jacobi form for $\mathrm{SL}_{2}(\mathbb{Z})$ of weight $1/2$ and index $1/2$. We work instead with a slightly modified version of this function. Readers familiar with [8] should be aware that $\widetilde{\Theta}(z;\tau)$ here is slightly different from the one in that paper. Namely, in terms of Dedekind’s eta-function $\eta(\tau):=q^{\frac{1}{24}}\prod_{n=1}^{\infty}(1-q^{n}),$ it will be convenient for us to employ
@@
The infinite product in (1.5) is taken over relatively prime pairs of integers $(m,n)$ instead of $i=1,2,\dots,$ as in (1.1). This modification does not lose any information. In fact, this reformulation will allow us to compute the values of the $\widehat{A}$-genus (and also $L$-genus) as quasimodular forms (see Theorem 1.4). Moreover, it is straightforward to reconstruct the original $\widehat{A}$-genus expressions (and also $L$-genus) using elementary properties of symmetric functions. The example after Theorem 1.4 illustrates this combinatorial and number theoretic procedure. Finally, we note that the dependence on $s>0$ in the index set $X_{\tau}(s)$ is required, as we view these series as analytic functions, and the introduction of $s$ guarantees convergence.
@@
As a consequence of both Theorems 1.1 and 1.2, we obtain quasimodular representations of the $\widehat{A}$-genus and $L$-genus. We now turn to the problem of converting these quasimodular forms into the original topological expressions. To make this precise, we make the important observation that these forms are given as traces of “partition Eisenstein series,” which are studied in [8, 12, 13]. To define them, recall that a partition of a non-negative integer $k$ (see [14] for background on partitions) is any nonincreasing sequence of positive integers
@@
$\widehat{A}_{3}(p_{1},p_{2},\dots)$ $=\frac{1}{967680}(-16p_{3}+44p_{1}p_{2}-31p_{1}^{3}),$
@@
(1.21) $\operatorname{Tr}_{k}^{(E)}(\phi;\tau):=\sum_{\lambda\vdash k}\phi(\lambda)E_{\lambda}(\tau),$
@@ Acknowledgements
The authors thank the two referees for their comments and suggestions. They also thank Jack Morava for helpful conversations and encouragement. They also thank John Shareshian and Richard Stanley for bringing our attention to Hitchin’s article [4], and they thank Wei-Lun Tsai for spotting typographical errors in an earlier version of this paper. The third author thanks the Thomas Jefferson Fund and the NSF (DMS-2002265 and DMS-2055118).
@@ 2.1. Jacobi forms and elliptic functions
Proof of Proposition 2.1. Claim (1) follows as an easy application of the Jacobi Triple Product formula (see Theorem 2.8 of [14]), which allows us to write
@@
The second claim follows immediately from (1) and (1.4). \ \ $\square$
@@
Proof. The $\sigma$-function has a $q$-series expansion (see Theorem I.6.6.4 of [17]) given by
@@
Thus $\sigma$ is also related to the modified theta function $\Theta(z;\tau)$, defined by (2.1), as claimed. \ $\square$
@@
Proof. The standard application of “Hecke’s trick” (for example, see p. 84 of [5]), to force convergence of the weight $2$ Eisenstein series, gives the formula
@@
The lemma follows since $\zeta(s)$ is continuous at $2$square
@@
In this section we prove Theorems 1.4 and 1.5, which express the quasimodular representations of the $\widehat{A}$-genus and $L$-genus as traces of partition Eisenstein series. To obtain these results, we make use of exponential generating functions that arose in the previous section. These generating functions can be reformulated as traces of partition Eisenstein series using special identities within the framework of Pòlya’s theory of cycle index polynomials.
@@ 3.1. Pòlya’s cycle index polynomials
The proof of claim (2) is a little more involved. Beginning with (1.17) and Theorem 1.4, we apply the correspondence $G_{2j}\longleftrightarrow B_{2j}s_{j}/2j$ as follows